Evaluate limπ‘₯β†’2 π‘₯^3βˆ’2π‘₯^2+4π‘₯βˆ’8/x^4βˆ’2π‘₯^3+π‘₯βˆ’2


1
Expert's answer
2021-06-07T14:54:45-0400

"lim_{𝑥→2} \frac{𝑥^3−2𝑥^2+4𝑥−8}{x^4−2𝑥^3+𝑥−2}\hspace{1cm}(\frac{0}{0}form)\\
\text{Apply L'Hospital Rule,}\\
=lim_{𝑥→2} \frac{3𝑥^2−4𝑥+4}{4x^3−6𝑥^2+1}\\
=\frac{8}{9}"